Mildred left camp 4 hours after margret did. they travelled in oposite directions. mildred drove at 30 Miles an hour, margeret at 45 miles an hour. how many hours after mildred left are the two girls 630 miles apart?
Found 2 solutions by rfer, checkley77:
Answer by rfer(16322)

M is 120 miles ahead
moving apart at 75 mph
630-120=510
510/75=6.8 hrs
4+6.8=10.8 hrs=10 hrs 48 min

Answer by checkley77(12844)

D=RT
630=30(T-4)+45T
630=30T-120+45T
75T=630+120
75T=750
T=750/75
T=10 HOURS AFTER MARGARET LEFT THEY WILL MEET.
PROOF:
630=30(10-4)+45*10
630=30*6+450
630=180+450
630=630